# Large Angular Momentum

### v3

@@ Abstract
Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere. A spin $j$ for a generic $j$ (a $12j+$ 1-state system) is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$. Here we study the state of a single angular momentum/spin in the limit, $j\to\infty$. The special class of states $|j,{\mathbf{n}}\rangle\in{\mathbf{CP}}^{2j}$, with spin oriented towards definite spatial directions ${\mathbf{n}}\in S^{2}$, i.e., $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, are found to behave as classical angular momenta, $j\,{\mathbf{n}}$, in this limit. Vice versa, general spin states in ${\mathbf{CP}}^{2j}$ do not become classical, even at large $j$. We discuss these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix. Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories for a particle carrying a large spin), and to make the widespread idea that large spins somehow become classical, more precise.
@@ 1 Introduction
It is a widely shared view that in the limit of large spin (angular momentum) a quantum-mechanical spin somehow becomes classical [1]-[21]. After all, the large spin limit ($j/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ ($j$  fixed) limit, and, in accordance with the general idea of the semi-classical limit in quantum mechanics, a large-angular-momentum state might be expected to behave as a classical angular momentum. But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs or if it does, exactly how. It is the aim of this note to help filling in this gap.
@@
The way how such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [22]. The magnetic moment of a macroscopic particle is the expectation value
@@
Typically, the spins are oriented along different directions, but the crucial point is that the component particles carrying them are bound in the macroscopic body in various atomic, molecular or crystalline structures. The wave function of particles in bound states does not split in sub-wavepackets, in contrast to that of an isolated spin $\tfrac{1}{2}$ atom, under an inhomogeneous magnetic field. The bound-state Hamiltonian does not allow it  [^1] This discussion is closely parallel to the question of quantum diffusion. Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials. This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies. As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is simply suppressed by mass.. What determines the motion of the CM of the body is the average magnetic moment, (1.3).
Logically, however, one cannot exclude special cases (e.g. a magnetized piece of metal) with all spins inside the body oriented in the same direction. One might wonder how the wave function of such a particle with spin $j\to\infty$ behaves, under an inhomogeneous magnetic field. The question is whether the three conditions recognized in [22] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reviewed in Appendix A here, are indeed sufficient. Or, is some extra condition, or a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig. 1 for a small spin, $j=13/2$) under an inhomogeneous magnetic field?
Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $J_{z}=m$, $-j\leq m\leq j$ ($m=-j+k$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots).
The aim of this work is to answer these questions. We shall first discuss in Sec. 2 some mathematical and physical aspects of the quantum states of a small and large angular momenta, and compare them with the properties of a classical angular momentum. Considerations of quantum fluctuations of various angular momentum states lead us to propose that a particular class of states, $|j,{\mathbf{n}}\rangle$,
@@
${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (1.5)
are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit. The states (1.4) are known in the literature as coherent spin states or Bloch states [5, 10, 6, 11, 7].
@@
In subsequent sections, we are going to examine these ideas through the analysis of the Stern-Gerlach processes (Sec. 3), the orbital angular momenta (Sec. 4), the angular-momentum addition rule (Sec. 5), and the rotation matrix (Sec. 6). Generic spin $j$ states far from (1.4) are shown to remain quantum mechanical even in the limit $j\to\infty$: the fate of these states will be discussed in Sec. 7. Conclusion and a few more general reflections are in Sec. 8.
@@ 2 Space of quantum spin states and classical angular momenta
The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the wave function
@@
$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}/\{0\}\,$ (2.2)
@@
$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (2.9)
@@
have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit. The states (2.15) are known as Bloch states or coherent spin states in the literature [5, 10, 6, 11, 7].
@@
$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}/{\{0\}}$ (2.21)
@@
where $\binom{n}{k}=n!/k!(n-k)!$ are the binomial coefficients. The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or coherent spin) states defined in [5, 10, 6, 11, 7], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
@@
Thus the stellar representation of $\mathbf{CP}^{n}$ points - the generic states of spin $j$ provides us with an appealing intuitive picture of quantum fluctuations. The $\mathbf{CP}^{n}$ points with randomly distributed stars, Fig. 3, represent spin states with large quantum fluctuations. The special states, (2.15), (2.25) - (2.27) with coincident stars, Fig. 4, instead, are those with the minimum fluctuations, and correspond to the Bloch or spin coherent state [5, 10, 6, 11, 7]. In the next four sections we shall illustrate how these latter states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
@@ 3 Stern-Gerlach experiments
In this section we compare the Stern-Gerlach experiment [27] for small and large spins, with special attention to the Bloch states, (2.15), (2.26), (2.27).
@@ 3.2 Large spin
Figure 5: The distribution $|c_{k}|^{2}$ in $k$ for possible values of $J_{z}=m$ ($m=-j+k$), for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak). The particle starts looking like classical.
Figure 6: The distribution $|c_{k}|^{2}$ ($J_{z}=m=-j+k$) as in Fig. 5 but for a spin $j=2\cdot 10^{5}$, for $\theta=\pi/2$ (center, orange peak) or $\theta=\pi/4$ (right, blue peak). This figure is drawn by using the approximation (3.7) and (3.8), rather than the exact formula, (2.26), (2.27).
@@ 7 General large-spin states
In Sec. 3 $\sim$ Sec. 6 we have seen how the spin coherent (Bloch) states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit. It is an entirely different story with a generic large spin $j$ state in ${\mathbf{CP}}^{2j}$,
@@
The status of general angular momentum states with strong fluctuations, (7.1), is similar to that of Einstein’s macroscopic split wave packet. In the large $j$ limit, a generic large spin state (represented by randomly distributed stars, Fig. 3), far from the Bloch state $|j,\mathbf{n}\rangle$, necessarily acquires a large space extension under an inhomogeneous magnetic field: a macroscopic quantum state. See Fig. 10. But a macroscopic pure quantum state is possible only at exceedingly low temperature close to $T=0$: otherwise, it is a mixture. See [22] for more discussions.
@@ 8 Conclusion
The observations made in this note render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one. In particular, we found the identification between the states with minimum fluctuations (2.15) - the coherent spin states - with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
@@
On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ remain quantum-mechanical even in the $j\to\infty$ limit. They can stay pure quantum states only at exceedingly low temperatures, and isolated. The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other. A macroscopic system is in a mixed state at finite temperatures [31, 32, 33],[22]. No macroscopically split pure wave packets in a SG setting survive decoherence. At the same time, as pointed out in [23], a macroscopic quantum state which becomes a mixture due to decoherence, remains such: a mixed quantum state. Decoherence in itself does not make the system considered classical, contrary to the statements often made in the literature.
The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a fruitful one both mathematically and physically, as seen in several occasions. From the mathematical point of view, the entire theory of angular momentum can be reconstructed this way [26], as we recalled already, and it helps to recover certain formulas easily. It is also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized (toy) model of a macroscopic body, made of many atoms and molecules (spins). Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$, Fig. 5, Fig. 6, can be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.  [^8] The suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is somewhat analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [34], to explain the general exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones. This being so, let us not forget that the deepest insight of Schrödinger in his book was that many biological processes such as mutation, reproduction, and inheritance, involve quantum mechanics in essential ways.
Nevertheless, logically our discussion is about the quantum and classical properties of a single large spin, and is not about the thermodynamical or statistical properties of many body systems, as in [6, 7], for instance. And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[21], the ideas around the proposal (2.16) exposed in Sec. 2, Sec. 3, Sec. 5, and Sec. 7 are, to the best of our knowledge, new.

### v5

@@ Abstract
The quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere. A spin $j$ for a generic $j$ (a $12j+$ 1-state system) is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$. Here we study the state of a single angular momentum/spin in the limit, $j\to\infty$. The special class of states $|j,{\mathbf{n}}\rangle\in{\mathbf{CP}}^{2j}$, with spin oriented towards definite spatial directions ${\mathbf{n}}\in S^{2}$, i.e., $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, are found to behave as classical angular momenta, $j\,{\mathbf{n}}$, in this limit. Vice versa, general spin states in ${\mathbf{CP}}^{2j}$ do not become classical, even at large $j$. We study these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix. Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories of a particle carrying a large spin in an inhomogeneous magnetic field), and to make the widespread idea that large spins somehow become classical, more precise.
@@ 1 Introduction
It is a widely shared view that in the limit of large spin (angular momentum) a quantum-mechanical spin somehow becomes classical [1]-[24]. After all, the large spin limit ($j/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ ($j$  fixed) limit, and, in accordance with the general idea of the semi-classical limit in quantum mechanics, a large-angular-momentum state might be expected to behave as a classical angular momentum. But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs, or if it does, exactly how. It is the aim of this note to help filling in this gap.
@@
The way such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [25]. The magnetic moment of a macroscopic particle is the expectation value
@@
Typically, the spins of the component atoms and molecules are oriented along different directions, but the crucial point is that they are bound in the macroscopic body in various atomic, molecular or crystalline structures. The wave function of particles in bound states does not split in sub-wavepackets, in contrast to that of an isolated spin $\tfrac{1}{2}$ atom, under an inhomogeneous magnetic field. The bound-state Hamiltonian does not allow it  [^1] This discussion is closely parallel to the question of quantum diffusion. Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials. This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies. As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is simply suppressed by mass.. What determines the motion of the CM of the body is the average magnetic moment, (1.3).
Logically, however, one cannot exclude special cases (e.g. a magnetized piece of metal) with all spins inside the body oriented in the same direction. One might wonder how the wave function of such a particle with spin $j\to\infty$ behaves, under an inhomogeneous magnetic field. The question is whether the three conditions recognized in [25] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reported here in Appendix A, are indeed sufficient. Or, is some extra condition, or perhaps a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig. 1 for a small spin, $j=13/2$) under an inhomogeneous magnetic field?
Figure 1: The distribution in $J_{z}=-j+k$ ($k=0,1,\ldots,2j$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots). The wave packet of such a particle spreads into many subwavepackets à la Stern-Gerlach under an inhomogeneous magnetic field with strong gradient in the ${\hat{z}}$ direction.
This is the central question we are going to investigate. We shall first discuss in Sec. 2 some mathematical and physical aspects of the quantum states of a small and large angular momenta, and compare them with the properties of a classical angular momentum. Considerations of quantum fluctuations of various angular momentum states lead us to propose that a particular class of states, $|j,{\mathbf{n}}\rangle$,
@@
or close to them, are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit.
@@
Lest the reader is led to a misunderstanding of the content of our work, let us make the following point clear. The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles is quite a fruitful one both from mathematical and physical point of view. From the mathematical point of view, the entire theory of angular momentum can indeed be reconstructed this way [29], and it will help to recover certain formulas for large angular momentum states easily. It is also useful from physics point of view, as such a system may be regarded as an idealized, toy model for a more realistic macroscopic body, made of many atoms and molecules (carrying spins and orbital angular momenta). And this picture helps to interprete some of our findings. (The states of type (1.4) are known, in such a context of many-body systems, as spin coherent states, Bloch states or Glauber coherent states, depending on the author [5, 11, 6, 12, 7, 9, 15, 19, 20]. They are all equivalent to the state of definite spin orientation (1.4), as far as the global spin property is concerned.)
Nevertheless, our discussion is, as will be clear from an attentive reading, entirely about the quantum or classical properties of a single large spin (or angular momentum). We are not concerned here with the thermodynamical or other physical properties of a many-body system. The main question is whether a single large spin system behaves classically or remains quantum mechanical, in the limit, $j/\hbar\to\infty$. As will be seen in the following, the answer turns out to be quite subtle and nontrivial.
In subsequent sections, we are going to examine these questions through the analysis of the Stern-Gerlach processes (Sec. 3), the orbital angular momenta (Sec. 4), the angular-momentum addition rule (Sec. 5), and the rotation matrix (Sec. 6). Generic spin $j$ states far from (1.4) are shown to remain quantum mechanical even in the limit $j\to\infty$: the fate of these states will be discussed in Sec. 7. Conclusion and a few more general reflections are in Sec. 8.
@@ 2 Space of quantum spin states and classical angular momenta
The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the vector
@@
$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}\,$ (2.2)
@@
$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}\;,$ (2.9)
@@
(already defined in (1.4) in the Introduction) have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit. The states (2.15) are known as Bloch states or coherent spin states in the literature [5, 11, 6, 12, 7].
@@
$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}$ (2.21)
@@
where $\binom{n}{k}=n!/k!(n-k)!$ are the binomial coefficients. The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or spin coherent) states considered in many-body physics context [5, 11, 6, 12, 7, 9, 15, 19, 20], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
@@
Thus the stellar representation of $\mathbf{CP}^{n}$ points - the generic states of spin $j$ provides us with an appealing intuitive picture of quantum fluctuations. The $\mathbf{CP}^{n}$ points with randomly distributed stars, Fig. 3, represent spin states with large quantum fluctuations. The special states, (2.15), (2.25) - (2.27) with coincident stars, Fig. 4, instead, are those with the minimum fluctuations, and correspond to the Bloch or spin coherent state [5, 11, 6, 12, 7]. In the next four sections we shall illustrate how these states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
@@ 3 Stern-Gerlach experiments
In this section we compare the Stern-Gerlach experiment [30] for small and large spins, with special attention to the states, (2.15), (2.26), (2.27).
@@ 3.2 Large spin
Figure 5: A distribution in possible values of $J_{z}=-j+k$, for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak). The graph represents the distribution $|c_{k}|^{2}$. The particle starts looking like classical.
Figure 6: The distribution in $J_{z}$ ($\equiv-j+k$) as in Fig. 5 but for a spin $j=2\cdot 10^{5}$, for $\theta=\pi/2$ (center, orange peak) or $\theta=\pi/4$ (right, blue peak). This figure is drawn by using the approximation (3.7) and (3.8), rather than the exact formula, (2.26), (2.27).
@@ 5 Addition of angular momenta
5.1 Spin-orbit interactions
As a possible variation of theme, and for completeness, let us consider a spin-orbit interaction of the form,
$H=A\,{\mathbf{L}}\cdot{\mathbf{S}}\;$ (5.18)
where $A$ is a constant, ${\mathbf{L}}$ and ${\mathbf{S}}$ are the orbital angular momentum and spin, respectively, of a given system. The problem is to understand whether and/or how such an interaction Hamiltonian reduces effectively to an interaction between two classical angular momenta, in the limit, $L/\hbar\to\infty$, $S/\hbar\to\infty$. By writing
$H=\frac{A}{2}\,({\mathbf{J}}^{2}-{\mathbf{L}}^{2}-{\mathbf{S}}^{2})\;,\qquad{\mathbf{J}}={\mathbf{L}}+{\mathbf{S}}\;,$ (5.19)
and assuming that ${\mathbf{L}}^{2}$ and ${\mathbf{S}}^{2}$ have fixed values $L(L+1)\hbar^{2}$ and $S(S+1)\hbar^{2}$ in the system under consideration, the problem reduces to that of angular momentum addition, already considered in this section. As such, an interaction of the form (5.18) would reduce to their classical counterpart, when the two states with angular momentum moduli $L$ and $S$ are the states of minimum fluctuations, such as $|\ell,\ell\rangle$ of (4.2) and $|j,{\mathbf{n}}\rangle$ of (2.15).
On the other hand, for generic atomic states such conditions are not expected to be satisfied, even in the limit of large atoms. Thus the discussion of this section is mostly irrelevant, and the standard, quantum mechanical analysis of spin-orbit interactions are required always.
@@ 7 General large-spin states
In Sec. 3 $\sim$ Sec. 6 we have seen how the particular spin states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit. It is an entirely different story with a generic large spin $j$ state in ${\mathbf{CP}}^{2j}$,
@@
The status of general angular momentum states with strong fluctuations, (7.1), is similar to that of Einstein’s macroscopic split wave packet. In the large $j$ limit, a generic large spin state (represented by randomly distributed stars, Fig. 3), far from the Bloch state $|j,\mathbf{n}\rangle$, necessarily acquires a large space extension under an inhomogeneous magnetic field: a macroscopic quantum state. See Fig. 10. But a macroscopic pure quantum state is possible only at exceedingly low temperature close to $T=0$: otherwise, it is a mixture.
To show quantitatively how a macroscopically split wave packet becomes a mixed state under an environment-induced decoherence, is a complex problem, depending on the details of the environment itself, the temperature, density, pressure, kinds of environment particles, and average density and momentum distribution, the mean de Broglie wave length, etc. It is beyond the scope of the present work to perform such a study. However, a qualitative discussion about how a macroscopically split wave packet becomes a mixture under an environment-induced decoherence [34, 35, 36], but not classical, has been given in [26], in the case of a spin $\frac{1}{2}$ particle.
@@ 8 Conclusion
The observations made in this work render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one. In particular, we found that the states with minimum fluctuations (2.15) and the states close to them, are to be identified with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
@@
On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ remain quantum-mechanical even in the $j\to\infty$ limit. They can stay pure quantum states only at exceedingly low temperatures, and isolated. The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other. A macroscopic system is in a mixed state at finite temperatures [34, 35, 36, 37],[25]. No macroscopically split pure wave packets in a SG setting survive decoherence. At the same time, as pointed out in [26], a macroscopic quantum state which becomes a mixture due to decoherence, remains such: a mixed quantum state. Decoherence in itself does not render the system involved classical, contrary to such an idea sometimes expressed in the literature [35, 37].
The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a useful one both mathematically and physically, as we saw in several occasions. The entire theory of angular momentum can indeed be reconstructed this way as shown by Schwinger [29], and it helps to understand certain formulas easily. It is also relevant from the physics point of view, as such a large angular momentum system may be regarded as an idealized toy-model version of a macroscopic body, made of many atoms and molecules (spins). Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$ we have observed, such as in Fig. 5, Fig. 6, can be interpreted as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.  [^8] The suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is somewhat analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [38], to explain the general exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones. This being so, let us not forget that the deepest insight of Schrödinger in his book was that many biological processes such as mutation, reproduction, and inheritance, involve quantum mechanics in essential ways.
Nevertheless, our discussion is really about the quantum and classical properties of a single large spin, and do not concern the thermodynamical, statistical or other physical properties of realistic many-body systems, as in [6, 7]. And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[24], our careful observations about how the particular class of large momentum (spin) states get classical in the limit, $j\to\infty$, independently of the temperature, whereas a generic large $j$ states remain quantum mechanical in the same limit, as illustrated in Sec. 2 - Sec. 5, and Sec. 7 are, to the best of our knowledge, new.